# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

### v1

@@ 2 Incremental Gradient Descent and Shuffled SGD
Since each $f_{t}$ is convex and $L$-smooth, we have for $t\in[T]$ (see (1.3)):
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}}{2L},$ (2.2)
$f_{t}({\bm{x}}_{k-1,t})-f_{t}({\bm{z}})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k-1,t}-{\bm{z}}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}}{2L}.$ (2.3)
On the other hand, letting $\Phi_{k-1,t}({\bm{x}}):=\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}\right\rangle+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{x}}\|^{2}$, we have $\nabla\Phi_{k-1,t}({\bm{x}}_{k-1,t+1})=\mathbf{0}$ by the update ${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla f_{t}({\bm{x}}_{k-1,t})$ in Algorithm 1. Observe that $\Phi_{k-1,t}$ is $\frac{1}{\eta_{k}}$-strong convex, and thus we also have
$\Phi_{k-1,t}({\bm{z}})\geq\Phi_{k-1,t}({\bm{x}}_{k-1,t+1})+\frac{1}{2\eta_{k}}\|{\bm{z}}-{\bm{x}}_{k-1,t+1}\|^{2}.$ (2.4)
Summing (2.2) and (2.3) and using (2.4), we conclude that for $t\in[T]$,
@@
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}$
[…]
$+\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t}\right\rangle}_{{\mathcal{T}}_{2}}+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}$
$-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}.$
@@
${\mathcal{T}}_{1}=\;$ $-\frac{1}{\eta_{k}}\sum_{t=1}^{T-1}\sum_{s=t+1}^{T}\left\langle{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t},{\bm{x}}_{k-1,s+1}-{\bm{x}}_{k-1,s}\right\rangle$
$-\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}$
@@
To make the first two terms on the right-hand side both nonpositive, we require
@@
Our next step is to specify our choice of the reference point ${\bm{z}}_{k}$ for each epoch $k$. In particular, we consider a sequence of points $\{{\bm{z}}_{k}\}_{k=-1}^{K-1}$ that is recursively defined as a convex combination of ${\bm{x}}_{k}$ and the previous reference point ${\bm{z}}_{k-1}$:
${\bm{z}}_{k}=(1-\lambda_{k}){\bm{x}}_{k}+\lambda_{k}{\bm{z}}_{k-1}$ (2.5)
for $k\geq 0$ with ${\bm{z}}_{-1}={\bm{x}}_{*}$ and $\lambda_{k}\in(0,1)$ to be set later.
Observe that ${\bm{z}}_{k}$ can also be written as a convex combination of the points $\{{\bm{x}}_{j}\}_{j=0}^{k}$ and ${\bm{x}}_{*}$ by unrolling the recursion, i.e.,
where $(1-\lambda_{k})+\prod_{i=0}^{k}\lambda_{i}+\sum_{j=0}^{k-1}\big{(}\prod_{i=j+1}^{k}\lambda_{i}\big{)}(1-\lambda_{j})=1$. If we set $\lambda_{k}=1$ for all $k$, then we have ${\bm{z}}_{k}={\bm{x}}_{*}$ and recover the bound $f({\bm{x}}_{k})-f({\bm{x}}_{*})$ in Lemma 2.1 which leads to the average iterate guarantee. For general $\{\lambda_{k}\}$, we prove the following lemma to relate the function value gap $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ to the optimality gap $f({\bm{x}}_{k})-f({\bm{x}}_{*})$.
Lemma 2.2.
Let ${\bm{z}}_{k}$ be defined by (2.5) for a given sequence of parameters $\lambda_{k}\in(0,1)$, where $k\geq 0$ and ${\bm{z}}_{-1}={\bm{x}}_{*}.$ Under 1, if there exists a sequence of nonnegative weights $w_{k}$ such that $\lambda_{k}w_{k}\leq w_{k-1}$ for $k\in[K-1]$, then for all $k\in[K]$:
  1. 1.
  2. 2.
@@
    Using (2.6) and the convexity of $f$, we have
@@
The role of the sequence of weights $\{w_{k}\}$ in Lemma 2.2 is to ensure that we can telescope the terms $\|{\bm{x}}_{k-1}-{\bm{z}}_{k-1}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}$ in (2.1). On the other hand, to succinctly see why such $\{{\bm{z}}_{k}\}$ could lead to the desired last iterate guarantees, we note that the second part of Lemma 2.2 indicates that $w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}$ intrinsically includes retraction terms of the optimality gaps at the previous iterates. Hence, we can deduct $w_{K-1}(f({\bm{x}}_{K})-f({\bm{x}}_{*}))$ from $\sum_{k=1}^{K}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}$ by properly choosing $\lambda_{k}$ and $w_{k}$ to cancel out the optimality gap terms at the intermediate iterates. In this case, the convergence rate for the last iterate is characterized by the growth rate of the sequence $\{w_{k}\}$.
Our choice of the reference points $\{{\bm{z}}_{k}\}$ is inspired by the recent works (Zamani and Glineur, 2023; Liu and Zhou, 2023) on last iterate guarantees for subgradient methods and SGD with $\lambda_{k}=\frac{w_{k-1}}{w_{k}}$. However, their proof techniques are not directly applicable to incremental methods, due to several technical obstacles including additional nontrivial error terms of the form $\frac{\alpha}{\beta}T\big{(}f({\bm{z}}_{k})-f({\bm{x}}_{*})\big{)}$ in (2.1) arising from the incremental gradient steps using reference points other than ${\bm{x}}_{*}$. Such error terms inherently deteriorate the growth rate of $\{w_{k}\}$ and could possibly lead to a worse last iterate rate compared to the rate on the average iterate. In the following theorem, we calibrate such degradation on the last iterate guarantees, and show that with a slightly smaller step size one can still achieve essentially the same rate as for the average iterate.
Theorem 2.3 (IGD).
Under Assumptions 1, 3, and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step size is fixed and satisfies $\eta_{k}=\eta\leq\frac{1}{\sqrt{\beta}TL}$, the output ${\bm{x}}_{K}$ of Algorithm 1 satisfies
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\mathrm{e}\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T\eta K^{\frac{1}{1+\alpha/\beta}}}.$ (2.7)
@@
Plugging ${\bm{z}}_{k-1}$ defined by (2.5) into (2.1) from Lemma 2.1 and using the inequality $\|\sum_{i=1}^{n}{\bm{x}}_{i}\|^{2}\leq n\sum_{i=1}^{n}\|{\bm{x}}_{i}\|^{2}$ to bound the term $\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}$, we obtain
@@
where $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$. Multiplying $\eta_{k}w_{k-1}$ on both sides with $w_{k}$ such that $\lambda_{k}w_{k}\leq w_{k-1}$ and noticing that $\|{\bm{x}}_{k-1}-{\bm{z}}_{k-1}\|^{2}\leq\lambda_{k-1}\|{\bm{x}}_{k-1}-{\bm{z}}_{k-2}\|^{2}$ by (2.5), we have
@@
$T\sum_{k=1}^{K}\eta_{k}\Big{[}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}-\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)}\Big{]}$ (2.8)
[…]
@@
Plugging back into (2.8), grouping the like terms, and choosing $\lambda_{0}=1$, we obtain
$T\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}$ (2.10)
@@
$\lambda_{k}w_{k}\leq\;$ $w_{k-1},\quad 0\leq k\leq K-1,$ (2.11)
$\eta_{k}w_{k-1}-\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\geq\;$ $1\displaystyle 0,\quad 1\leq k\leq K-$ 1 . (2.12)
Noticing that (2.12) is equivalent to $\lambda_{k}\geq 1-\frac{\eta_{k}w_{k-1}}{\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}\sum_{j=k+1}^{K}\eta_{j}}$, to have both inequalities satisfied at the same time, it suffices that
@@
To maximize the growth rate of $\{w_{k}\}$, we let $w_{k}=\frac{\eta_{k}+(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}{(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}$. Without loss of generality, we take $w_{-1}=w_{0}=\prod_{k=1}^{K-1}\frac{(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}{\eta_{k}+(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}$, and thus $w_{K-1}=1$. Hence, dividing both sides of (2.10) by $T\eta_{K}w_{K-1}$ and choosing the constant step size $\eta_{k}\equiv\eta$ for all $k\in[K]$, we obtain
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\eta^{2}T^{2}\sigma_{*}^{2}L\sum_{k=1}^{K}w_{k-1}+\frac{w_{-1}}{2\eta T}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}.$ (2.13)
@@
Observe that the error term $f({\bm{z}}_{k-1})-f({\bm{x}}_{*})$ in (2.1) plays the role of slowing the last iterate rate, as is calibrated by its coefficient $\alpha/\beta$ in (2.7). To remedy such degradation compared to the average iterate rate (Mishchenko et al., 2020; Cai et al., 2023a), one natural thought is to make $\alpha/\beta$ sufficiently small. In particular, we choose $\alpha/\beta=1/\log K$ and show that the last iterate rate nearly matches the best known rate on the average iterate, with the trade-off of requiring order-$\frac{1}{\sqrt{\log K}}$ smaller step sizes in comparison with Mishchenko et al. (2020); Cai et al. (2023a). This translates into the oracle complexity that is larger by at most a $\sqrt{\log(1/\epsilon)}$ factor. For most cases of interest, this quantity can be treated as a constant; in particular, for $\epsilon=10^{-8},$ $\sqrt{\log(1/\epsilon)}\approx 4.$
On the other hand, we note that (2.11) and (2.12) hold trivially with $\lambda_{k}\equiv 1$ and constant $w_{k}$, and then (2.10) directly implies the average iterate guarantee, as a sanity check. Also, instead of saturating (2.11) and (2.12) to obtain the last iterate guarantee, one can deduce the convergence rate on the increasing weighted averaging which places more weight on later iterates, as formalized in the following corollary.
Corollary 2.4 (Increasing Weighted Averaging).
Under Assumptions 1, 3, and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the constant step size $\eta\leq\frac{1}{\sqrt{\beta}TL}$, then for any constant $c\in(0,1]$ and increasing sequence $\{w_{k}\}_{k=0}^{K-1}$ with $w_{k}=\frac{(1+\frac{\alpha}{\beta})(K-k)+1-c}{(1+\frac{\alpha}{\beta})(K-k)}w_{k-1}$lgorithm 1 outputs $\hat{\bm{x}}_{K}=\sum_{k=1}^{K}\frac{w_{k-1}{\bm{x}}_{k}}{\sum_{k=1}^{K}w_{k-1}}$ satisfying
@@
We follow the proof of Theorem 2.3 up to (2.10) with constant step size $\eta$, then we instead take $\lambda_{k}=w_{k-1}/w_{k}$ and $w_{k}=\frac{(1+\frac{\alpha}{\beta})(K-k)+1-c}{(1+\frac{\alpha}{\beta})(K-k)}w_{k-1}$ to obtain
@@
Since $f$ is convex, we have $f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{\sum_{k=1}^{K}w_{k-1}(f({\bm{x}}_{k})-f({\bm{x}}_{*}))}{\sum_{k=1}^{K}w_{k-1}}$ where $\hat{\bm{x}}_{K}=\sum_{k=1}^{K}\frac{w_{k-1}{\bm{x}}_{k}}{\sum_{k=1}^{K}w_{k-1}}$ is the increasing weighted averaging of $\{{\bm{x}}_{k}\}_{k=1}^{K}$, thus (cf. (2.13))
@@
$TK={\mathcal{O}}\Big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c^{3/2}\epsilon^{3/2}}\Big{)}.$
We remark that increasing weighted averaging shaves off the (at most) square-root-log term appearing in the last iterate rate above and recovers the best known rate for the average iterate (Mishchenko et al., 2020; Cai et al., 2023a). The parameters $\alpha,\beta$ are included in the weights $w_{k}$ controlling the growth rate of the increasing sequence $\{w_{k}\}$. In particular, when $\alpha/\beta\rightarrow\infty$, increasing weighted averaging reduces to the uniform weighted average.
In the final part of this section, we extend our analysis to handle the case with possible random permutations on the task ordering of each epoch, showing order-$\sqrt{T}$ improvements in complexity if involving randomness. We consider two main permutation strategies of particular interests in the literature on shuffled SGD: (i) random reshuffling (RR): randomly generate permutations at the beginning of each epoch; (ii) shuffle-once (SO): generate a single random permutation at the beginning and use it in all epochs. We state the convergence results with permutations in the following corollary. Our last iterate guarantee nearly matches the best known average iterate convergence rate for shuffled SGD (Mishchenko et al., 2020; Nguyen et al., 2021; Cai et al., 2023a) and the lower bound results on the RR scheme (Cha et al., 2023), with a slightly (order-$(1/\sqrt{\log K})$) smaller step size.
Corollary 2.5 (Shuffled SGD (RR/SO)).
Under Assumptions 13 and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the constant step size $\eta\leq\frac{1}{\sqrt{\beta}TL}$, the output ${\bm{x}}_{K}$ of Algorithm 1 with uniformly random (SO/RR) shuffling satisfies
$\mathbb{E}[f({\bm{x}}_{K})-f({\bm{x}}_{*})]\leq\;$ $\mathrm{e}\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T\eta K^{\frac{1}{1+\alpha/\beta}}}.$

### v2

@@ Appendix A Omitted Proofs From Section 2
Since each $f_{t}$ is convex and $L$-smooth, we have for $t\in[T]$ (see Eq. (1.3)):
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t}\right\rangle-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2},$ (A.1)
$f_{t}({\bm{x}}_{k-1,t})-f_{t}({\bm{z}})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k-1,t}-{\bm{z}}\right\rangle-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}.$ (A.2)
On the other hand, letting $\Phi_{k-1,t}({\bm{x}}):=\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}\right\rangle+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{x}}\|^{2}$, we have $\nabla\Phi_{k-1,t}({\bm{x}}_{k-1,t+1})=\mathbf{0}$ by the update ${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla f_{t}({\bm{x}}_{k-1,t})$ in Alg. 1. Observe that $\Phi_{k-1,t}$ is $\frac{1}{\eta_{k}}$-strong convex, and thus we also have
$\Phi_{k-1,t}({\bm{z}})\geq\Phi_{k-1,t}({\bm{x}}_{k-1,t+1})+\frac{1}{2\eta_{k}}\|{\bm{z}}-{\bm{x}}_{k-1,t+1}\|^{2}.$ (A.3)
Summing Eq. (A.1) and (A.2) and using Eq. (A.3), we conclude that for $t\in[T]$,
@@
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}}_{{\mathcal{T}}_{1}}$
$+\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t}\right\rangle}_{{\mathcal{T}}_{2}}$
$-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}$
@@
${\mathcal{T}}_{1}=\;$ $-\frac{1}{\eta_{k}}\sum_{t=1}^{T-1}\sum_{s=t+1}^{T}\left\langle{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t},{\bm{x}}_{k-1,s+1}-{\bm{x}}_{k-1,s}\right\rangle-\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}$
@@
To make the first two terms on the right-hand side both nonpositive, we choose
@@
See 2.2
@@
    Using Eq. (2.2) and the convexity of $f$, we have
@@
See 2.3
@@
Plugging ${\bm{z}}_{k-1}$ defined by Eq. (2.1) into Lemma 2.1 and using the inequality $\|\sum_{i=1}^{n}{\bm{x}}_{i}\|^{2}\leq n\sum_{i=1}^{n}\|{\bm{x}}_{i}\|^{2}$ to bound the term $\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}$, we obtain
@@
where $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$. Multiplying $\eta_{k}w_{k-1}$ on both sides with $w_{k}$ such that $\lambda_{k}w_{k}\leq w_{k-1}$ and noticing that $\|{\bm{x}}_{k-1}-{\bm{z}}_{k-1}\|^{2}\leq\lambda_{k-1}\|{\bm{x}}_{k-1}-{\bm{z}}_{k-2}\|^{2}$ by Eq. (2.1), we have
@@
$T\sum_{k=1}^{K}\eta_{k}\Big{[}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}-\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)}\Big{]}$ (A.4)
$\leq$ $T^{3}\sigma_{*}^{2}L\sum_{k=1}^{K}\eta_{k}^{3}w_{k-1}+\frac{w_{-1}}{2}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{\alpha}{\beta}T\sum_{k=1}^{K}\eta_{k}\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)},$
@@
Plugging back into Eq. (A.4), grouping the like terms, and choosing $\lambda_{0}=1$, we obtain
$T\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}$ (A.6)
@@
$\lambda_{k}w_{k}\leq\;$ $w_{k-1},\quad 0\leq k\leq K-1,$ (A.7)
$\eta_{k}w_{k-1}-\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\geq\;$ $1\displaystyle 0,\quad 1\leq k\leq K-$ 1 . (A.8)
Noticing that Eq. A.8 is equivalent to $\lambda_{k}\geq 1-\frac{\eta_{k}w_{k-1}}{\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}\sum_{j=k+1}^{K}\eta_{j}}$, to have both inequalities satisfied at the same time, it suffices that
@@
To maximize the growth rate of $\{w_{k}\}$, we let $w_{k}=\frac{\eta_{k}+(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}{(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}$. Without loss of generality, we take $w_{-1}=w_{0}=\prod_{k=1}^{K-1}\frac{(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}{\eta_{k}+(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}$, and thus $w_{K-1}=1$. Hence, dividing both sides of Eq. (A.6) by $T\eta_{K}w_{K-1}$ and choosing the constant step size $\eta_{k}\equiv\eta$ for all $k\in[K]$, we obtain
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\eta^{2}T^{2}\sigma_{*}^{2}L\sum_{k=1}^{K}w_{k-1}+\frac{w_{-1}}{2\eta T}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}.$ (A.9)
@@
See 2.4
@@
We follow the proof of Theorem 2.3 up to Eq. (A.6) with constant step size $\eta$, then we instead take $\lambda_{k}=w_{k-1}/w_{k}$ and $w_{k}=\frac{(1+\frac{\alpha}{\beta})(K-k)+1-c}{(1+\frac{\alpha}{\beta})(K-k)}w_{k-1}$ to obtain
@@
Since $f$ is convex, we have $f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{\sum_{k=1}^{K}w_{k-1}(f({\bm{x}}_{k})-f({\bm{x}}_{*}))}{\sum_{k=1}^{K}w_{k-1}}$ where $\hat{\bm{x}}_{K}=\sum_{k=1}^{K}\frac{w_{k-1}{\bm{x}}_{k}}{\sum_{k=1}^{K}w_{k-1}}$ is the increasing weighted averaging of $\{{\bm{x}}_{k}\}_{k=1}^{K}$, thus (cf. Eq. (A.9))
@@
$TK={\mathcal{O}}\Big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c^{3/2}\epsilon^{3/2}}\Big{)},$
thus finishing the proof. ∎
See 2.5